# On sufficient conditions for the transitivity of homeomorphisms

### v3

@@ 7.3.2. Generic elements of $\mathcal{WIP}_{h}$
()($\subseteq$)( ⊆ ) Take $f\in\mathcal{R}\,\cap\,\mathcal{WIP}_{h}$, that is, $f\in\mathcal{KS}$ and $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$. By Corollary 7.3, all periodic points of $f$ are hyperbolic and have the same index. Thus, $f\in\mathcal{PCI}$. By Proposition 7.7, $f$ is $\Omega-$stable. Since, in addition, $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$, from Theorem C we deduce that $f$ is chain-transitive. Hence, $f\in\mathcal{R}\,\cap\,\mathcal{PCI}\,\cap\,\mathcal{CT}$.
()superset-of-or-equals($\supseteq$)( ⊇ ) Suppose that $f\in\mathcal{R}\,\cap\,\mathcal{PCI}\cap\mathcal{CT}$. By Proposition 7.7, $f$ is $\Omega-$ stable. Thus, $f$ is chain-transitive. Therefore, by Theorem C, $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$. ∎
@@
$(7)\Rightarrow(4)$ Assume that $f$ has the specification property. By [53, Lemma 2.1], we know that $f$ has the weak $su-$intersecting property in $\operatorname{Per}_{h}(f)$.
@@
$(9)\Rightarrow(4)$ Suppose that $f$ has the gluing orbit property. By Lemma 6.2, $f$ has the barycenter property in $A$ for every nonempty $A$ of $M$. Since $\mathcal{R}\subseteq\mathcal{KS}$, we have $\emptyset\neq\operatorname{Per}(f)=\operatorname{Per}_{h}(f)$. Thus, $f$ has the barycenter property in $\operatorname{Per}_{h}(f)$, and so it has the weak $su-$intersecting property in $\operatorname{Per}(f)$. ∎
@@ 7.4. Application 4
[…]
@@
When the dimension of the manifold is equal to $3$ or $4$ it is known that a $C^{1}-$robustly transitive diffeomorphism may not be an Anosov diffeomorphism [38, 54]. On the other hand, in [8], A. Arbieto and R. Ribeiro showed that the $C^{1}$ interior of the set of flows with the asymptotic average shadowing property (or with the average shadowing property) on a $3-$dimensional smooth closed manifold is formed by topologically transitive Anosov flows.
@@ 7.5. Application 5
    $f$ has the weak $su-$intersecting property in $\operatorname{Per}_{h}(f)$.
@@
If $f\in\mathcal{KS}_{m}$ and $f$ has the weak $su-$intersecting property in $\operatorname{Per}(f)$, then all the periodic points have the same index.
@@
$(1)\Rightarrow(2)$ Suppose that $f$ has the barycenter property in $\operatorname{Per}_{h}(f)$. Since $f\in\mathcal{KS}_{m}$, one has $\operatorname{Per}_{h}(f)=\operatorname{Per}(f)$, and so $f$ satisfies the weak $su-$intersecting property in $\operatorname{Per}_{h}(f)$.
$(2)\Rightarrow(3)$ Assume that $f$ has the weak $su-$intersecting property in $\operatorname{Per}_{h}(f)$.
@@ Acknowledgments
L.S. was partially supported by FAPERJ-Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro Projects APQ1-E-26/211.690/2021 SEI-260003/015270/2021 and JCNE-E-26/200.271/2023 SEI-260003/000640/2023, by Coordenação de Aperfeiçoamento de Pessoal de Nível Superior CAPES - Finance Code 001 and PROEXT-PG project Dynamic Women - Dinâmicas, CNPq-Brazil (grant Projeto Universal 404943/2023-3). MC has been partially supported by Portuguese public funds through FCT - Fundação para a Ciência e a Tecnologia, I.P., under the projects with references UIDB/00144/2020, UIDP/00144/2020 and PTDC/MAT-PUR/4048/2021.

### v4

@@ 7.3.2. Generic elements of $\mathcal{WIP}_{h}$
$(\subseteq)$ Take $f\in\mathcal{R}\,\cap\,\mathcal{WIP}_{h}$, that is, $f\in\mathcal{KS}$ and $f$ has the weak susu-intersecting property in $\operatorname{Per}(f)$. By Corollary 7.3, all periodic points of $f$ are hyperbolic and have the same index. Thus, $f\in\mathcal{PCI}$. By Proposition 7.7, $f$ is $\Omega$-stable. Since, in addition, $f$ has the weak susu-intersecting property in $\operatorname{Per}(f)$, from Theorem C we deduce that $f$ is chain-transitive. Hence, $f\in\mathcal{R}\,\cap\,\mathcal{PCI}\,\cap\,\mathcal{CT}$.
$(\supseteq)$ Suppose that $f\in\mathcal{R}\,\cap\,\mathcal{PCI}\cap\mathcal{CT}$. By Proposition 7.7, $f$ is $\Omega$- stable. Thus, $f$ is chain-transitive. Therefore, by Theorem C, $f$ has the weak susu-intersecting property in $\operatorname{Per}(f)$. ∎
@@
$(7)\Rightarrow(4)$ Assume that $f$ has the specification property. By [53, Lemma 2.1], we know that $f$ has the weak susu-intersecting property in $\operatorname{Per}_{h}(f)$.
@@
$(9)\Rightarrow(4)$ Suppose that $f$ has the gluing orbit property. By Lemma 6.2, $f$ has the barycenter property in $A$ for every nonempty $A$ of $M$. Since $\mathcal{R}\subseteq\mathcal{KS}$, we have $\emptyset\neq\operatorname{Per}(f)=\operatorname{Per}_{h}(f)$. Thus, $f$ has the barycenter property in $\operatorname{Per}_{h}(f)$, and so it has the weak susu-intersecting property in $\operatorname{Per}(f)$. ∎
@@ 7.4. Application 4
$\text{topological transitivity}\quad\Rightarrow\quad\text{chain-transitivity}.$
@@
When the dimension of the manifold is equal to $3$ or $4$ it is known that a $C^{1}$-robustly transitive diffeomorphism may not be an Anosov diffeomorphism [38, 54]. On the other hand, in [8], A. Arbieto and R. Ribeiro showed that the $C^{1}$ interior of the set of flows with the asymptotic average shadowing property (or with the average shadowing property) on a 33-dimensional smooth closed manifold is formed by topologically transitive Anosov flows.
@@ 7.5. Application 5
    $f$ has the weak susu-intersecting property in $\operatorname{Per}_{h}(f)$.
@@
If $f\in\mathcal{KS}_{m}$ and $f$ has the weak susu-intersecting property in $\operatorname{Per}(f)$, then all the periodic points have the same index.
@@
$(1)\Rightarrow(2)$ Suppose that $f$ has the barycenter property in $\operatorname{Per}_{h}(f)$. Since $f\in\mathcal{KS}_{m}$, one has $\operatorname{Per}_{h}(f)=\operatorname{Per}(f)$, and so $f$ satisfies the weak susu-intersecting property in $\operatorname{Per}_{h}(f)$.
$(2)\Rightarrow(3)$ Assume that $f$ has the weak susu-intersecting property in $\operatorname{Per}_{h}(f)$.
@@ Acknowledgments
L.S. was partially supported by FAPERJ-Fundação Carlos Chagas Filho de Amparo à Pesquisa do Estado do Rio de Janeiro Projects APQ1-E-26/211.690/2021 SEI-260003/015270/2021 and JCNE-E-26/200.271/2023 SEI-260003/000640/2023; CAPES-Coordenação de Aperfeiçoamento de Pessoal de Nível Superior - Finance Code 001 and PROEXT-PG project Dynamic Women - Dinâmicas; CNPq-Brazil (grant Projeto Universal 404943/2023-3). MC has been partially supported by Portuguese public funds through FCT - Fundação para a Ciência e a Tecnologia, I.P., under the projects with references UIDB/00144/2020, UIDP/00144/2020 and PTDC/MAT-PUR/4048/2021.